# thanks: This letter has been submitted in conjunction with its companion manuscript Zhang et al. (2024)

### v1

@@ Abstract
A novel goodness-of-fit strategy is introduced for testing models for angular power spectra characterized by unknown parameters. Using this strategy, it is possible to assess the validity of such models without specifying the distribution of the estimators of the angular power spectrum being used. This holds under general conditions, ensuring the applicability of the method across diverse scenarios. Moreover, the proposed solution overcomes the need for case-by-case simulations when testing different models – leading to notable computational advantages.
@@ I Motivation and goal
The angular power spectrum characterizes the power distribution of quantities on a sphere across angular scales. Understanding its features and assessing the validity of the corresponding theoretical models is critical in studying a variety of fundamental phenomena in different scientific fields. For instance, in geodesy, it is used to characterize the properties of the Earth’s sea surface topography Hwang (1993) as well as the gravitational and magnetic fields of Earth and other planets Albertella et al. (1999); Simons and Dahlen (2006); Lesur (2006); Livermore et al. (2024). It is also used in geophysics to infer mechanical properties of the lithosphere Wieczorek (2015), in atmospheric science to examine surface temperature anomalies Evans (1998); Sneppen (2022), and in medical imaging applications Maniar and Mitra (2005); Mitra and Maniar (2006). Of particular interest to us is its relevance in cosmology and astrophysics. For instance, in the context of the cosmic microwave background (CMB), angular power spectra of both CMB temperature and polarization have enabled precise measurements of essential cosmological parameters, such as the energy density of dark matter and dark energy  Collaboration et al. (2020). Galaxy surveys, like the Sloan Digital Sky Survey (SDSS) Ahumada et al. (2020), have provided insights into the galaxy distribution across the sky. The angular power spectrum derived from these distributions has shed light on the evolution of visible matter in the universe Hayes and Brunner (2012); Yang et al. (2023). Similarly, weak gravitational lensing surveys, such as the Dark Energy Survey (DES), have generated weak lensing convergence sky-maps and their angular power spectrum describes the mass distribution in the universe Jeffrey et al. (2021); Amon et al. (2022). Measurements of the stochastic gravitational-wave (GW) background energy density across the sky and its angular power spectrum could be used to identify the astrophysical or cosmological mechanism(s) that generated the background Geller et al. (2018); Cusin et al. (2018, 2019, 2017); Jenkins et al. (2018); Jenkins and Sakellariadou (2018), albeit to date only upper limit constraints have been made Abbott et al. (2021). Additionally, cross-correlations of these GW sky maps and their electromagnetic counterparts (galaxy counts, weak lensing, etc.) have been proposed as auxiliary astrophysical probes Cusin et al. (2018, 2019, 2017), and first measurements have been made Yang et al. (2023). This trend is anticipated to persist and intensify with the release of new telescopes and detectors in the upcoming decade, with a tenfold (or greater) surge in the acquired data volumes.
@@
Let $T(\widehat{\Omega},x)$ be a sky map with $\widehat{\Omega}$ being the sky direction and $x$ denoting other observables, such as redshift or frequency bin of the observed light or GWs. For example, $T(\widehat{\Omega},x)$ may represent the number of galaxies observed over different pixels and at a given redshift, or the CMB temperature measured across the sky and at a given CMB frequency bin. To ease the exposition we focus on the situation in which $x\in[\mathcal{L},\mathcal{U}]\subseteq\mathbb{R}$. Extensions to multidimensional observables are not difficult and are discussed at the end of this letter.
Consider the decomposition of $T(\widehat{\Omega},x)$ given by
$T(\widehat{\Omega},x)$ $=\sum_{\ell=0}^{\infty}\sum_{m=-\ell}^{\ell}a_{\ell m}(x)Y_{\ell m}(\widehat{\Omega})$ (1)
with $Y_{\ell m}(\widehat{\Omega})$ being the spherical harmonics, that is, complex-valued basis functions defined on a sphere, and
@@
are the corresponding coefficients. The angular power spectrum is defined as
$C_{\ell}(x)$ $=$ $\frac{1}{2\ell+1}\sum_{m=-\ell}^{\ell}|a_{\ell m}(x)|^{2}.$ (2)
@@
In practical applications, the sky map $T(\widehat{\Omega},x)$ is often directly measured in the pixel basis and can then be converted into estimates $\widehat{a}_{\ell m}(x)$ using (1). In some cases, the available data covers only a limited portion of the sky, hence resulting in large experimental uncertainties for $\widehat{a}_{\ell m}(x)$ estimators at low $\ell$’s. In other cases, such as when measuring the anisotropic stochastic GW background, it is possible to directly estimate the $a_{\ell m}(x)$ coefficients (and the corresponding covariance matrix) by cross-correlating the time-series strain data from two geographically separated GW detectors Allen and Romano (1999); Thrane et al. (2009); Abbott et al. (2021). In any case, by replacing $a_{\ell m}(x)$ with $\widehat{a}_{\ell m}(x)$ in (2), and for any value of $x$ fixed, we can obtain estimates $\widehat{C}_{\ell}(x)$, $\ell=1,\dots,L$ with $L$ being the point of truncation of the expansion in (1). Hereafter assume the latter to be unbiased estimates of $C_{\ell}(x)$ for all $\ell=1,\dots,L$ (see Zhang et al. (2024) for details on the construction of unbiased estimators).
@@
Let $\bm{C}^{M}(x,\bm{\theta})$ be a theoretical model with components $C^{M}_{\ell}(x,\bm{\theta})$ aiming to predict $C_{\ell}(x)$ in (2). Denote with $\widehat{\bm{C}}(x)$ the vector with components $\widehat{C}_{\ell}(x)$. Once an estimate of the unknown parameter $\bm{\theta}\in\mathbb{R}^{p}$ is obtained, our goal is to assess the validity of $\bm{C}^{M}(x,\bm{\theta})$ based on the information carried by $\widehat{\bm{C}}(x)$ and its covariance matrix $\bm{\Sigma}(x)$.
@@
Various methods for estimating $\bm{C}^{M}(x,\bm{\theta})$ have been proposed and often rely on the assumption that, for each $x$ fixed, $\widehat{\bm{C}}(x)$ is a Gaussian random vector Netterfield et al. (2002); Knox and Page (2000); Piacentini et al. (2006); Jungman et al. (1996); Knox (1995). However, while the estimates $\widehat{a}_{\ell m}(x)$ are usually obtained by averaging large amounts of data and are asymptotically distributed as a complex multivariate Gaussian, the $\widehat{C}_{\ell}(x)$ are averages of the squares of $\widehat{a}_{\ell m}(x)$. It follows that their distribution is that of a generalized $\chi^{2}$ whose likelihood function does not have a closed-form expression Zhang et al. (2024) and, in general, cannot be considered approximately Gaussian. Solutions that aim to account for non-Gaussianity typically involve likelihoods, which are combinations of a Gaussian and offset log-normal distribution Bennett et al. (2003, 2013). However, those have also been shown to be unsatisfactory in some settings Sun et al. (2013). Unfortunately, when the distribution of $\widehat{\bm{C}}(x)$ is not adequately modeled, constructing a goodness-of-fit test for $\bm{C}^{M}(x,\bm{\theta})$ is particularly challenging. Nonetheless, such a test is unavoidable to ensure the reliability of the resulting estimates of $\bm{\theta}$ and subsequent conclusions.
Motivated by this challenge, we discuss a distribution-free approach to testing theoretical models for angular power spectra. As detailed in Sec. II, one can easily access a suitable estimation procedure of the parameters by reformulating the problem within the context of regression. The latter allows us to employ recent advancements in the theory of empirical processes (Khmaladze, 2021) and adapts them to our context to formally test the model $\bm{C}^{M}(x,\bm{\theta})$ without knowing the distribution of $\widehat{\bm{C}}(x)$. Moreover, this approach avoids the use of case-by-case simulations to derive the null distribution of the test statistics employed to perform goodness-of-fit, an especially desirable feature when $\bm{C}^{M}(x,\bm{\theta})$ has a complicated structure, as in the case of astrophysical anisotropic stochastic GW backgrounds or when cross-correlating different sky maps such as GW and galaxy counts sky maps Cusin et al. (2018, 2019); Jenkins et al. (2018); Cusin et al. (2017).
@@
This letter provides a self-contained exposition of the method in general settings. In a companion paper (Zhang et al., 2024), we present all of the technical details for the main results, discuss their extensions, and illustrate the implementation of the methodology to study the anisotropic stochastic gravitational-wave background. Finally, while the proposed procedure is motivated by the need for distribution-free inference for autocorrelation angular power spectra, the statistical tools described here also apply to the analysis of cross-correlation angular power spectra.
II Estimation via Least Squares
@@
Let $\mathcal{L}=x_{0}<x_{1}<\dots<x_{n}=\mathcal{U}$ be the values of the observable $x$ considered and denote with $\widehat{\bm{C}}$ the vector of components $\widehat{C}_{\ell}(x_{i})$, with $\ell=1,\dots,L$ and $i=1,\dots,n$, i.e., if $\widehat{\bm{C}}^{T}(x_{i})=[\widehat{C}_{1}(x_{i}),\ldots,\widehat{C}_{L}(x_{i})]$, then $\widehat{\bm{C}}$ is a vector of dimension $N=Ln$ and defined as
$\ \widehat{\bm{C}}=[\widehat{\bm{C}}^{T}(x_{1}),\ldots,\widehat{\bm{C}}^{T}(x_{n})]^{T}.$
@@
For example, $\widehat{\bm{C}}(x_{i})$ may represent the estimated angular power spectrum up to the $L$-th multipole for the GW energy density at the GW frequency $x_{i}$. Hence, $\widehat{\bm{C}}$ collects the estimated angular power spectra at frequencies $x_{1},\dots,x_{n}$. The vector $\bm{C}^{M}(\bm{\theta})$ of elements $C^{M}_{\ell}(x_{i},\bm{\theta})$ can be specified in a similar manner.
@@
Let $\bm{\Sigma}$ be the covariance matrix of $\widehat{\bm{C}}$. We assume the random vectors $\widehat{\bm{C}}(x_{i})$ at different $x_{i}$ values to be independent so that, for all $\ell,\ell^{\prime}=1,\dots,L$ and $i\neq i^{\prime}$,
$\text{cov}\bigl{(}\widehat{C}_{\ell}(x_{i}),\widehat{C}_{\ell^{\prime}}(x_{i^{\prime}})\bigl{)}=0.$
@@
Therefore, $\bm{\Sigma}$ is a block diagonal matrix with blocks $\bm{\Sigma}(x_{i})$. We further assume that $\bm{\Sigma}$ is either known exactly, well approximated, or estimated by means of a consistent estimator Zhang et al. (2024). As an example, in the context of the anisotropic stochastic GW background, the covariance matrix for the $\widehat{a}_{\ell m}(x_{i})$ estimators can be estimated directly from the time-series strain data for each frequency bin $x_{i}$ Allen and Romano (1999); Thrane et al. (2009); Abbott et al. (2021) and, due to the properties of the Fourier transform, the estimators in different frequency bins are independent to a very good approximation. This covariance matrix can then be used to construct $\bm{\Sigma}$, which, in turn, is block-diagonal, with each block corresponding to a different GW frequency $x_{i}$.
@@
Estimating the astrophysical parameter $\bm{\theta}$ characterizing $\bm{C}^{M}(\bm{\theta})$ can be accomplished using least squares estimation. That is, by solving
$\begin{split}\widehat{\bm{\theta}}=\arg\min_{\bm{\theta}}(\widehat{\bm{C}}-\bm{C}^{M}(\bm{\theta}))^{T}\Sigma^{-1}(\widehat{\bm{C}}-\bm{C}^{M}(\bm{\theta})).\end{split}$ (3)
@@
Notice that the optimization problem (3), known in the statistic literature as ‘Generalized Nonlinear Least Squares’ (e.g., Davidson and MacKinnon, 2004, Ch. 6-7) is equivalent to estimating $\bm{\theta}$ under the assumption that the likelihood of $\widehat{\bm{C}}$ is approximately Gaussian (e.g., Ade et al., 2014; Hinshaw et al., 2003; Sun et al., 2013). However, the rationale behind (3) is simply that of minimizing the sum of the squares of the decorrelated residuals collected in the vector
$\widehat{\bm{\epsilon}}=\bm{\Sigma}^{-1/2}(\widehat{\bm{C}}-\bm{C}^{M}(\widehat{\bm{\theta}}))$
where $\bm{\Sigma}^{-1/2}$ is the inverse square-root matrix of $\bm{\Sigma}$ obtained, for example, via Jordan or Schur decomposition (e.g., Higham, 2008). It is also worth emphasizing that $\widehat{\bm{\theta}}$ can be shown to be a consistent estimator  [^1] A consistent estimator is an estimator whose mean converges to the true value of the parameter being estimated and its variance converges to zero for sufficiently large samples. of $\theta$ on the basis of the classical theory of $M$-estimation (Huber, 1964) and does not require Gaussianity of $\widehat{\bm{C}}$  [^2] When Gaussianity holds, $\widehat{\bm{\theta}}$ in (3) is not only a consistent estimator of $\bm{\theta}$ but also has minimum variance.. Hence, what is relevant to ensure the consistency of $\widehat{\bm{\theta}}$ as an estimator of $\bm{\theta}$ is not the Gaussianity of $\widehat{\bm{C}}$ but rather the validity of the postulated model $\bm{C}^{M}(\bm{\theta})$. Therefore, our goal is to test the hypothesis:
$H_{0}:\langle\widehat{\bm{C}}\rangle=\bm{C}^{M}(\bm{\theta}).$ (4)

### v5

@@ Abstract
A novel goodness-of-fit strategy is introduced for testing models of angular power spectra with unknown parameters. Using this strategy, it is possible to assess the validity of such models without specifying the distribution of the angular power spectrum estimators. This holds under general conditions, ensuring the method’s applicability in diverse applications. Moreover, the proposed solution overcomes the need for case-by-case simulations when testing different models, leading to notable computational advantages.
@@ I Motivation and goal
The angular power spectrum characterizes the power distribution of quantities on a sphere across angular scales. Understanding its features and assessing the validity of the corresponding theoretical models is critical in studying a variety of fundamental phenomena in different scientific fields. For instance, in geodesy it is used to characterize the properties of the Earth’s sea surface topography Hwang (1993) as well as the gravitational and magnetic fields of Earth and other planets Albertella et al. (1999); Simons and Dahlen (2006); Lesur (2006); Livermore et al. (2024). It is also used in geophysics to infer mechanical properties of the lithosphere Wieczorek (2015), in atmospheric science to examine surface temperature anomalies Evans (1998); Sneppen (2022), and in medical imaging applications Maniar and Mitra (2005); Mitra and Maniar (2006). Of particular interest to us is its relevance in cosmology and astrophysics. For instance, in the context of the cosmic microwave background (CMB), angular power spectra of both CMB temperature and polarization have enabled precise measurements of essential cosmological parameters, such as the energy density of dark matter and dark energy  Collaboration et al. (2020). Galaxy surveys, like the Sloan Digital Sky Survey (SDSS) Ahumada et al. (2020), have provided insights into the galaxy distribution across the sky. The angular power spectrum derived from these distributions has shed light on the evolution of visible matter in the universe Hayes and Brunner (2012); Yang et al. (2023). Similarly, weak gravitational lensing surveys, such as the Dark Energy Survey (DES), have generated weak lensing convergence sky-maps and their angular power spectrum describes the mass distribution in the universe Jeffrey et al. (2021); Amon et al. (2022). Measurements of the stochastic gravitational-wave (GW) background energy density across the sky and its angular power spectrum could be used to identify the astrophysical or cosmological mechanism(s) that generated it Geller et al. (2018); Cusin et al. (2018, 2019, 2017); Jenkins et al. (2018); Jenkins and Sakellariadou (2018), albeit to date only upper limit constraints have been made Thrane et al. (2009); Abbott et al. (2017, 2019, 2021, 2022); Agarwal et al. (2022); Renzini and Contaldi (2019a, b). Additionally, cross-correlations of these GW background sky maps and their electromagnetic counterparts (galaxy counts, weak lensing, etc.) have been proposed as auxiliary astrophysical probes Cusin et al. (2018, 2019, 2017), and first measurements have been made Yang et al. (2023)  [^1] While there are studies considering cross-correlations between electromagnetic signals and distributions of individually resolvable GW sources Calore et al. (2020); Zazzera et al. (2024); Ghosh et al. (2025); Namikawa et al. (2016); Oguri (2016); Raccanelli et al. (2016); Scelfo et al. (2018), we refer specifically to a GW background signal.. This trend is anticipated to persist and intensify with the release of new telescopes and detectors in the upcoming decade, with a tenfold (or greater) surge in the acquired data volumes.
@@
Let $T(\widehat{\Omega},\mathbf{x})$ be a sky map with $\widehat{\Omega}$ being the sky direction and $\mathbf{x}$ denoting other independent variables, such as redshift or frequency bin of the observed light or GWs. For example, $T(\widehat{\Omega},\mathbf{x})$ may represent the number of galaxies observed over different pixels and at a given redshift, or the CMB temperature measured across the sky and at a given CMB frequency bin. Note that $\mathbf{x}$ can be multidimensional, i.e., $\mathbf{x}\in\mathcal{X}\subseteq\mathbb{R}^{D}$.
Consider the decomposition of $T(\widehat{\Omega},\mathbf{x})$ given by
@@
[…]
$a_{\ell m}(\mathbf{x})=\int T(\widehat{\Omega},\mathbf{x})Y_{\ell m}(\widehat{\Omega})d\widehat{\Omega}$. (2)
The angular power spectrum is defined as
$C_{\ell}(\mathbf{x})$ $=$ $\frac{1}{2\ell+1}\sum_{m=-\ell}^{\ell}|a_{\ell m}(\mathbf{x})|^{2}.$ (3)
@@
In practical applications, the sky map $T(\widehat{\Omega},\mathbf{x})$ is often directly measured in the pixel basis and can then be converted into estimates $\widehat{a}_{\ell m}(\mathbf{x})$ using (1). In some cases, the available data covers only a limited portion of the sky, hence resulting in large experimental uncertainties for $\widehat{a}_{\ell m}(\mathbf{x})$ estimators at low $\ell$’s. In other cases, such as when measuring the anisotropic stochastic GW background, it is possible to directly estimate the $a_{\ell m}(\mathbf{x})$ coefficients (and the corresponding covariance matrix) by cross-correlating the time-series strain data from two geographically separated GW detectors Allen and Romano (1999); Thrane et al. (2009); Abbott et al. (2021). In any case, by replacing $a_{\ell m}(\mathbf{x})$ with $\widehat{a}_{\ell m}(\mathbf{x})$ in (3), and for any value of $\mathbf{x}$ fixed, we can obtain estimates $\widehat{C}_{\ell}(\mathbf{x})$, $\ell=1,\dots,L$ with $L$ being the point of truncation of the expansion in (1). Hereafter assume the latter to be unbiased estimates of $C_{\ell}(\mathbf{x})$ for all $\ell=1,\dots,L$ (see Zhang et al. (2025) for details on the construction of unbiased estimators). Note that while $\ell=0$ can be included, we choose not to do so since often one is interested only in fluctuations of $T(\widehat{\Omega},\mathbf{x})$ around its mean, implying $C_{0}(\mathbf{x})=0$.
@@
Let $\mathbf{C}^{M}(\mathbf{x},\mathbf{\theta})$ be a theoretical model with components $C^{M}_{\ell}(\mathbf{x},\mathbf{\theta})$ aiming to predict $C_{\ell}(\mathbf{x})$ in (3). Denote with $\widehat{\mathbf{C}}(\mathbf{x})$ the vector with components $\widehat{C}_{\ell}(\mathbf{x})$. Once an estimate of the unknown parameter $\mathbf{\theta}\in\mathbb{R}^{p}$ is obtained, our goal is to assess the validity of $\mathbf{C}^{M}(\mathbf{x},\mathbf{\theta})$ based on the information carried by $\widehat{\mathbf{C}}(\mathbf{x})$ and its covariance matrix $\bm{\Sigma}(\mathbf{x})$.
@@
Various methods for estimating $\mathbf{C}^{M}(\mathbf{x},\mathbf{\theta})$ have been proposed and often rely on the assumption that, for each $\mathbf{x}$ fixed, $\widehat{\mathbf{C}}(\mathbf{x})$ is a Gaussian random vector Netterfield et al. (2002); Knox and Page (2000); Piacentini et al. (2006); Jungman et al. (1996); Knox (1995). However, while the estimates $\widehat{a}_{\ell m}(\mathbf{x})$ are usually obtained by averaging large amounts of data and are asymptotically distributed as a complex multivariate Gaussian, the $\widehat{C}_{\ell}(\mathbf{x})$ are averages of the squares of $\widehat{a}_{\ell m}(\mathbf{x})$. It follows that their distribution is that of a generalized $\chi^{2}$ whose likelihood function does not have a closed-form expression Zhang et al. (2025) and, in general, cannot be considered approximately Gaussian. Solutions that aim to account for non-Gaussianity typically involve likelihoods that are combinations of a Gaussian and offset log-normal distribution Bennett et al. (2003, 2013). However, those have also been shown to be unsatisfactory in some settings Sun et al. (2013). Unfortunately, when the distribution of $\widehat{\mathbf{C}}(\mathbf{x})$ is not adequately modeled, constructing a goodness-of-fit test for $\mathbf{C}^{M}(\mathbf{x},\mathbf{\theta})$ is particularly challenging. Nonetheless, such a test is required to ensure the reliability of the resulting estimates of $\mathbf{\theta}$ and subsequent conclusions.
Motivated by this challenge, we discuss a novel distribution-free approach to testing theoretical models for angular power spectra. Traditionally, distribution-free goodness-of-fit tests are defined as tests that rely on test statistics with an asymptotic null distribution independent of the model or distribution being tested. Here, however, it acquires a broader meaning: the null distribution of the test statistics used does not depend on the model. Additionally, to derive this distribution, one does not need to know the distribution of the random variables whose mean is being tested, in our case, $\widehat{\mathbf{C}}(x)$. This is made possible by reformulating the estimation problem of $\mathbf{C}^{M}(\mathbf{x},\mathbf{\theta})$ as a regression problem and implementing the testing procedure proposed by Khmaladze (2021). This letter is the first work applying such a technique to enable distribution-free testing of models for the angular power spectrum.
A significant computational advantage is that a distribution-free testing solution avoids the need for case-by-case simulations when testing different models. It is, therefore, especially desirable when testing models with a complicated structure, as in the case of astrophysical anisotropic stochastic GW backgrounds or when cross-correlating different sky maps such as galaxy count maps and GW background sky maps Cusin et al. (2018, 2019); Jenkins et al. (2018); Cusin et al. (2017).
@@
This letter provides a self-contained exposition of the method in general settings. In a companion paper (Zhang et al., 2025), we present all of the technical details for the main results, discuss their extensions, and illustrate the implementation of the methodology to study the anisotropic stochastic gravitational-wave background. Finally, while the proposed procedure is motivated by the need for distribution-free inference for autocorrelation angular power spectra, the statistical tools described here also apply to the analysis of cross-correlation angular power spectra.
II Estimation via Generalized Least Squares
Let $\mathbf{x}_{1},\dots,\mathbf{x}_{n}$ be the vectors collecting the values of the independent variables. We assume that such values are linearly ordered so that, if $\mathcal{X}\subseteq\mathbb{R}$, $\mathbf{x}_{1}\leq\mathbf{x}_{2}\leq\dots\leq\mathbf{x}_{n}$. When $\mathcal{X}\subseteq\mathbb{R}^{D}$, we assume that the $D$-dimensional vectors $\mathbf{x}_{i}=(x_{i1},\dots,x_{iD})$ are ordered by increasing one dimension at a time. For example, if $D=2$ and $n=n_{1}\cdot n_{2}$, then:
$\begin{split}\mathbf{x}_{1}&=(x_{11},x_{12}),\mathbf{x}_{2}=(x_{21},x_{12}),\dots,\mathbf{x}_{n_{1}}=(x_{n_{1}1},x_{12}),\\ &\mathbf{x}_{n_{1}+1}=(x_{11},x_{22}),\dots,\mathbf{x}_{2n_{1}}=(x_{n_{1}1},x_{22}),\text{ etc.}\end{split}$ (4)
@@ II Estimation via Generalized Least Squares
Denote with $\widehat{\mathbf{C}}$ the vector of components $\widehat{C}_{\ell}(\mathbf{x}_{i})$, with $\ell=1,\dots,L$ and $i=1,\dots,n$, i.e., if $\widehat{\mathbf{C}}^{T}(\mathbf{x}_{i})=[\widehat{C}_{1}(\mathbf{x}_{i}),\ldots,\widehat{C}_{L}(\mathbf{x}_{i})]$, then $\widehat{\mathbf{C}}$ is a vector of dimension $N=L\cdot n$ and defined as
[…]
@@
For example, $\widehat{\mathbf{C}}(\mathbf{x}_{i})$ may represent the estimated angular power spectrum up to the $L$-th multipole for the GW energy density at the GW frequency $\mathbf{x}_{i}$. Hence, $\widehat{\mathbf{C}}$ collects the estimated angular power spectra at frequencies $\mathbf{x}_{1},\dots,\mathbf{x}_{n}$. The vector $\mathbf{C}^{M}(\mathbf{\theta})$ of elements $C^{M}_{\ell}(\mathbf{x}_{i},\mathbf{\theta})$ are specified similarly.
@@
Let $\bm{\Sigma}$ be the covariance matrix of $\widehat{\mathbf{C}}$. We assume the random vectors $\widehat{\mathbf{C}}(\mathbf{x}_{i})$ at different $\mathbf{x}_{i}$ values to be independent so that, for all $\ell,\ell^{\prime}=1,\dots,L$ and $i\neq i^{\prime}$,
$\text{cov}\bigl(\widehat{C}_{\ell}(\mathbf{x}_{i}),\widehat{C}_{\ell^{\prime}}(\mathbf{x}_{i^{\prime}})\bigl)=0.$ (6)
@@
Therefore, $\bm{\Sigma}$ is a block diagonal matrix with blocks $\bm{\Sigma}(\mathbf{x}_{i})$. We further assume that $\bm{\Sigma}$ is either known exactly, well approximated, or estimated by means of a consistent estimator Zhang et al. (2025). As an example, in the context of the anisotropic stochastic GW background, the covariance matrix for the $\widehat{a}_{\ell m}(\mathbf{x}_{i})$ estimators can be estimated directly from the time-series strain data for each frequency bin $\mathbf{x}_{i}$ Allen and Romano (1999); Thrane et al. (2009); Abbott et al. (2021) and the estimators in different frequency bins are independent to a very good approximation (Thrane et al., 2009; Allen and Romano, 1999). This covariance matrix can then be used to construct the block diagonal matrix $\bm{\Sigma}$, with each block corresponding to a different GW frequency $\mathbf{x}_{i}$.
@@
Estimating the astrophysical parameter $\mathbf{\theta}$ characterizing $\mathbf{C}^{M}(\mathbf{\theta})$ can be accomplished using least squares estimation. That is, by solving
$\begin{split}\widehat{\mathbf{\theta}}=\arg\min_{\mathbf{\theta}}(\widehat{\mathbf{C}}-\mathbf{C}^{M}(\mathbf{\theta}))^{T}\bm{\Sigma}^{-1}(\widehat{\mathbf{C}}-\mathbf{C}^{M}(\mathbf{\theta})).\end{split}$ (7)
@@
Notice that the optimization problem in (7), known in the statistical literature as ‘Generalized Nonlinear Least Squares’ (e.g., Davidson and MacKinnon, 2004, Ch. 6-7) is equivalent to using maximum likelihood for estimating $\mathbf{\theta}$ when the likelihood of $\widehat{\mathbf{C}}$ is approximately Gaussian (e.g., Ade et al., 2014; Hinshaw et al., 2003; Sun et al., 2013). However, the rationale behind (7) is simply that of minimizing the sum of the squares of the decorrelated residuals collected in the vector
$)\widehat{\mathbf{\epsilon}}=[\widehat{\epsilon}_{1},\dots,\widehat{\epsilon}_{N}]^{T}=\bm{\Sigma}^{-1/2}(\widehat{\mathbf{C}}-\mathbf{C}^{M}(\widehat{\mathbf{\theta}}$)) (8)
where $\bm{\Sigma}^{-1/2}$ is the inverse square-root matrix of $\bm{\Sigma}$ obtained, for example, via Jordan or Schur decomposition (e.g., Higham, 2008). The estimator $\widehat{\mathbf{\theta}}$ is consistent based on the classical theory of $M$-estimation (Huber, 1964). That is, regardless of the distribution of $\widehat{\mathbf{C}}$, as $N\rightarrow\infty$, $\widehat{\mathbf{\theta}}$ converges to the true value of $\theta$ when $\mathbf{C}^{M}(\theta)$ is correctly specified. Such a property is also necessary to enable the validity of the testing approach described in III (Cf. Zhang et al. (2025) for the technical details).
It follows that what is relevant to ensure the consistency of $\widehat{\mathbf{\theta}}$ as an estimator of $\mathbf{\theta}$ is not the Gaussianity of $\widehat{\mathbf{C}}$ but rather the validity of the postulated model $\mathbf{C}^{M}(\mathbf{\theta})$. Therefore, our goal is to test the hypothesis:
$H_{0}:\langle\widehat{\mathbf{C}}\rangle=\mathbf{C}^{M}(\mathbf{\theta}).$ (9)